OpenStudy (anonymous):

Multiply. [(x^2)/(x-5)]*[(x^2-7x+10)/(x^2+3x)

OpenStudy (jhannybean):

\[\large \frac{x^2}{x-5} \times \frac{x^2-7x+10}{x^2+3x}\] First factor out the GCF. \[\large \frac {x^2}{x-5}\times \frac{(x-5)(x-2)}{x(x+3)} \] And then cancel out the like terms.\[\large \frac {\cancel{x^2}x}{\cancel{x-5}} \times \frac{\cancel{(x-5)}(x-2)}{\cancel{x}(x+3)}\]

OpenStudy (jhannybean):

Rewrite your result.\[\large \frac{x(x-2)}{x+3}\]

OpenStudy (jhannybean):

Does that make sense?

OpenStudy (anonymous):

Yes

OpenStudy (jhannybean):

Awesome :D

OpenStudy (anonymous):

Thank you :3

OpenStudy (jhannybean):

noo problem ^_____^